Please help me with this problem. I have asked for hrlp before but could not find the answer to it. Thank you for your help. This problem is from a homework sheet.
Write the series in summation notation. Use the index i and let i begin at 1 in each summation.
3.) 9 + 14 + 19 + 24 + 29

Found 2 solutions by josmiceli, stanbon:
Answer by josmiceli(6799)
Answer by stanbon(48580)

Write the series in summation notation. Use the index i and let i begin at 1 in each summation.
9 + 14 + 19 + 24 + 29
a(n) = 9 + 5(n-1)
a(n) = 9 + 5n-5
a(n) = 4+5n
-----------------
Summation Notation:
Sum from i=1 to n of (4+5n)
========================
Cheers,
Stan H.